# XLIP: Cross-modal Attention Masked Modelling for Medical Language-Image Pre-Training

### v2

@@ 3.3 Entity-driven Masked Language Modeling Module
Figure 4: Illustration of entity-driven mask generation, which combines the CausalMask with the mask guided by the medical NER to generate the final result.
@@
$\mathbb{M}_{\text{EntMLM}}=\text{RandomSubSet}(M_{e},\lambda_{3})\cap\text{CausalMask}.$ (10)
@@
$\text{MLM}_{\text{input}}=\mathbb{M}_{\text{EntMLM}}\cdot E_{\text{report}}.$ (11)
Models CheXpert PadChest-Seen 8 Pneumonia
AUC F1 ACC AUC F1 ACC AUC F1 ACC
ConVIRT [48] 52.10 35.61 57.43 74.31 23.58 80.12 79.21 55.67 75.08
GLORIA [11] 54.84 37.86 60.70 74.56 24.02 80.75 70.37 48.19 70.54
BioViL [3] 60.01 42.10 66.13 71.17 20.75 79.58 84.12 54.49 74.43
BioViL-T [3] 70.93 47.21 69.96 76.17 27.41 85.32 86.03 62.56 80.04
CheXzero [36] 87.90 61.90 81.17 76.75 28.68 87.37 85.13 61.49 78.34
MedKlip [41] 89.97 63.67 84.32 84.97 35.17 90.68 86.94 63.42 80.02
XLIP 90.53 67.05 86.04 85.79 33.50 91.19 86.16 62.57 80.46
+ Ensemble 91.45 69.18 86.28 86.31 35.61 92.16 86.95 63.30 80.51
Table 1: Zero-shot performance of various pre-training models on three datasets.
3.4 Objective Functions
3.4.1 Objective function for image-report alignment.
Inspired by [21], our approach focuses on aligning image and text features prior to their fusion by the multimodal decoder. This pre-alignment strategy simplifies the process of cross-modal learning for the multimodal decoder, facilitating a more effective integration of visual and textual information. We follow [32] and compute the objective alignment function $Loss_{\text{align}}$ by exploiting the fine-grained correspondences between $E_{\text{img}}$ and $E_{\text{report}}$, formulated as:
@@ 3.4 Objective Functions
3.4.2 Objective function for image reconstruction.
The image decoder, inspired by [10], consists of an 8-layer transformer, which is essential for pre-training image reconstruction. This component integrates encoded visible patches and mask tokens to reconstruct the image. We defined this decoder as $D_{\text{MIM}}(\cdot)$. Reconstruction accuracy is quantified by the Mean Squared Error loss computed solely on masked patches, shown as:
$Loss_{\text{MIM}}=\left\|y^{\text{MIM}},x\right\|^{2},~{}\text{where}~{}y^{\text{MIM}}=D_{\text{MIM}}(\text{MIM}_{\text{input}}).$ (12)
3.4.3 Objective function for report reconstruction.
The text decoder in our model, following [47], features a 3-layer transformer. The multi-modal decoder combines visual and textual data, enhancing the representations. It uses a combination of causal and disease-guided masks and integrates visual encoder outputs via alignment from contrastive learning and the cross-attention mechanism.
@@
$Loss_{\text{MLM}}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}|y_{<t},x),~{}\text{where}~{}y_{t}\in\text{MLM}_{\text{input}}.$ (13)
3.4.4 Final objective function.
In the medical field, the scenario of unpaired data is common and significant, mainly due to data access limitations, ethical and privacy considerations, and the diversity and complexity of medical data. These factors make it difficult to obtain perfectly paired medical data in practical applications and also highlight the need to develop advanced algorithms capable of handling such data. XLIP Integrating prompt-driven attention and image self-attention facilitates the generation of masks for image data, which lacks corresponding reports. This approach improves the training process, supports compatibility with unpaired image data, and ensures that even data solely involved in the MIM task contributes to enhanced model representation learning. Our whole training process mixes paired data and unpaired data, thus the final objective function consists of these two parts as well.
@@
$Loss_{\text{pair}}=Loss_{\text{align}}+Loss_{\text{MIM}}+Loss_{\text{MLM}}.$ (14)
@@
$Loss_{\text{unpair}}=Loss_{\text{MIM}}.$ (15)
@@
$Loss_{\text{final}}=Loss_{\text{pair}}+Loss_{\text{unpair}}.$ (16)
4 Experiments
4.1 Dataset
@@
The experimental data is divided into pre-training data and downstream task data. The pre-training data includes MIMIC [16] data that can complete multi-modal tasks and Padchest [5] data used as unpaired image-only data. The downstream task data includes binary classification datasets CovidX [31] and Pneumonia [1], as well as multi-label classification of 14 types of diseases datasets CheXpert [13] and Xray14 [38]. Please refer to Appendix 6.1 for details.
4.2 Implementation Details
@@
We initialize the image and text encoders with official clip pre-training weights. The experiments are conducted using a single NVIDIA A100. SGD is employed as the network optimizer with a momentum of 0.9 and an initial learning rate of 5e-5. The input resolution is set to 224 × 224, and the batch size is 64. For a warm-up, we first optimize the masked modeling objectives for 15,000 iterations and then jointly train all objectives end-to-end. After pre-training, we keep only the image and text encoders for downstream tasks. Using the CLIP [32] approach, zero-shot classification is done by matching the image and category name features from both encoders. Meanwhile, the image encoder can be used independently for fine-tuning classification tasks. About evaluation metrics, following previous works [36, 42], we adopt the area under the ROC curve (AUC), ACC, and F1 scores, which are standard metrics to evaluate classification tasks.
5 Results
5.1 Zero-shot Classification
@@
We compare our XLIP with different pre-training methods, including ConVIRT [48], GLORIA [11], BioViL [3], BioViL-T[3], CheXzero [36], and MedKlip [41], for zero-shot classification. The results in Table 1 and Table 2 show that our XLIP exhibits superior zero-shot performance across all these scenarios on three datasets, even for the unseen class recognition on the PadChest dataset. This may be because it is difficult for existing models to understand the feature information contained in the whole image and report, while the proposed AttMIM and EntMLM can utilize the rich multimodal information of the medical data itself to help our XLIP to understand the more critical medical knowledge, and even understand the complex medical features that have not been mentioned in the report of the training set.
@@
Specifically, in Table 1, compared to the second-best MedKlip, the AUC scores of our XLIP have increased as follows: from 89.97 to 90.53 on the CheXpert dataset, from 84.97 to 85.79 on the PadChest-seen dataset, and from 85.94 to ours 86.16 on the RSNA Pneumonia dataset. Meanwhile, we used the ensemble learning method to combine the weights obtained from 6 different groups of parameters, which can improve the performance further on most datasets. In Table 2, we comparatively analyze the zero-shot performance of XLIP on the PadChest dataset for unseen classes. Only 8 categories in PadChest strictly match the visible categories in the MIMIC data, so we only report the test results of 8 visible categories in Table 1. There are 12 categories that either cannot be strictly matched, but cannot be counted as unseen classes, or do not appear in the test set, so we report at most 173 unseen classes. Our analysis of the 10 and 20 most numerous of these unseen classes reveals that the zero-shot performance of XLIP has a 4% to 5% lift.
@@
CheXzero [36] 59.75 63.11 65.76
MedKlip [41] 60.07 53.73 64.16
XLIP 65.09 67.14 70.00
Table 2: Comparison of zero-shot AUC scores on the PadChest dataset. Unseen 10 represents the 10 categories with the highest number of unseen categories.
Figure 5: Performance comparison under different labeled data for fine-tuning. Unpaired data like PadChest can work with pair data to improve image encoder performance on the fine-tuning task.
5.2 Fine Tuning
Existing multi-modal frameworks are constrained as they require data that include both text and images, which limits the potential for the image encoder to leverage image-only data to improve its representation capabilities. To address this, we introduce a training approach designed to work effectively with unpaired, single-modal data. This new method allows the image encoder to better use image-only data, enhancing its representation power and overall performance. Our experiments involved two datasets, where we used 1%, 10%, and 100% of the data for fine-tuning, aligning with the methods in previous works [41]. As shown in Figure 5, our model significantly outperforms existing models in terms of AUC scores across all datasets. Meanwhile, the performance of the model was further improved by incorporating image data from PadChest, demonstrating that compatibility with unpaired data is vital for representation learning.
Figure 6: Comprehensive evaluation of AUC and F1 under different mask ratio settings of AttMIM.
Figure 7: Comprehensive evaluation of AUC and F1 under different mask ratio settings of EntMLM.
5.3 Ablation Study
@@
5.3.1 Improvements from AttMIM
As shown in Table 4, when we randomly mask 75% of the features according to the setting of MIM in MAE [10] and directly splice this module with CLIP and initialize it with the weights of MAE decoder, we can see that the model performance improves by 0.42% of AUC score when we train it directly. However, CLIP and MIM rely on different training strategies, with the latter tending to require smaller learning rates and longer training epochs. To harmonize the two, we first train the image encoder via the MIM reconstruction task for 15,000 iterations, called Warm Up process. With the Warm Up, the model further improves the AUC by 1.1%. Meanwhile, we comparatively validate varying single strategies with mask ratios of 0.75, and the AttMIM strategy with the fusion of the three strategies. Experimental results show that three attention-based mask strategies can help the model understand key features better than random. As a fusion of the three strategies, AttMIM can more comprehensively improve related capabilities.
@@ 5.3.1 Improvements from AttMIM
Table 3: Performance of different image features masking strategies on the new CheXpert validation set.
@@
Table 4: Performance of different text features masking strategies on the new CheXpert validation set.
5.3.2 Improvements from EntMLM
@@
6 Conclusion
In this paper, we propose XLIP, an advanced medical VLP framework that aims to overcome major challenges in medical image and text analysis by introducing an innovative cross-modal attention masking modeling mechanism. By fusing AttMIM and EntMLM, XLIP can accurately capture key pathological features in medical data and improve the model’s ability to understand complex medical scenes. In addition, we also use disease prompts to make the model compatible with unpaired data. Through pre-training on two large-scale medical datasets and validation on multiple downstream tasks, XLIP demonstrates its superior performance, not only achieving state-of-the-art results in the zero-shot setting but also in the fine-tuning setting. This result proves the practicality and effectiveness of XLIP in deep learning applications in the medical field.

### v3

@@ 3.3. Entity-driven Masked Language Modeling Module
Figure 5. Illustration of entity-driven mask generation, which combines the CausalMask with the mask guided by the medical NER to generate the final result.
@@
(10) $\mathbb{M}_{\text{EntMLM}}=\text{RandomSubSet}(M_{e},\lambda_{3})\cap\text{CausalMask}.$
@@
(11) $\text{MLM}_{\text{input}}=\mathbb{M}_{\text{EntMLM}}\cdot E_{\text{report}}.$
3.4. Objective Functions
3.4.1. Objective function for image-report alignment.
Inspired by (Li et al., 2021), our approach focuses on aligning image and text features prior to their fusion by the multimodal decoder. This pre-alignment strategy simplifies the process of cross-modal learning for the multimodal decoder, facilitating a more effective integration of visual and textual information. We follow (Radford et al., 2021) and compute the objective alignment function $Loss_{\text{align}}$ by exploiting the fine-grained correspondences between $E_{\text{img}}$ and $E_{\text{report}}$, formulated as:
@@ 3.4. Objective Functions
3.4.2. Objective function for image reconstruction.
The image decoder, inspired by (He et al., 2022), consists of an 8-layer transformer, which is essential for pre-training image reconstruction. This component integrates encoded visible patches and mask tokens to reconstruct the image. We defined this decoder as $D_{\text{MIM}}(\cdot)$. Reconstruction accuracy is quantified by the Mean Squared Error loss computed solely on masked patches, shown as:
(12) $Loss_{\text{MIM}}=\left\|y^{\text{MIM}},x\right\|^{2},~{}\text{where}~{}y^{\text{MIM}}=D_{\text{MIM}}(\text{MIM}_{\text{input}}).$
3.4.3. Objective function for report reconstruction.
The text decoder in our model, following (Yu et al., 2022), features a 3-layer transformer. The multi-modal decoder combines visual and textual data, enhancing the representations. It uses a combination of causal and disease-guided masks and integrates visual encoder outputs via alignment from contrastive learning and the cross-attention mechanism.
@@
(13) $Loss_{\text{MLM}}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}|y_{<t},x),~{}\text{where}~{}y_{t}\in\text{MLM}_{\text{input}}.$
3.4.4. Final objective function.
In the medical field, the scenario of unpaired data is common and significant, mainly due to data access limitations, ethical and privacy considerations, and the diversity and complexity of medical data. These factors make it difficult to obtain perfectly paired medical data in practical applications and also highlight the need to develop advanced algorithms capable of handling such data. MMCLIP Integrating prompt-driven attention and image self-attention facilitates the generation of masks for image data, which lacks corresponding reports. This approach improves the training process, supports compatibility with unpaired image data, and ensures that even data solely involved in the MIM task contributes to enhanced model representation learning. Our whole training process mixes paired data and unpaired data, thus the final objective function consists of these two parts as well.
@@
(14) $Loss_{\text{pair}}=Loss_{\text{align}}+Loss_{\text{MIM}}+Loss_{\text{MLM}}.$
@@
(15) $Loss_{\text{unpair}}=Loss_{\text{MIM}}.$
@@
(16) $Loss_{\text{final}}=Loss_{\text{pair}}+Loss_{\text{unpair}}.$
4. Experiments
4.1. Dataset
@@
The experimental data is divided into pre-training data and downstream task data. The pre-training data includes MIMIC (Johnson et al., 2019) data that can complete multi-modal tasks and Padchest (Bustos et al., 2020) data used as unpaired image-only data. The downstream task data includes binary classification datasets CovidX (Pavlova et al., 2022) and Pneumonia (Anouk Stein, 2018), as well as multi-label classification of 14 types of diseases datasets CheXpert (Irvin et al., 2019) and Xray14 (Wang et al., 2017). Please refer to Appendix for details.
4.2. Implementation Details
@@
We initialize the image and text encoders with official clip pre-training weights. The experiments are conducted using a single NVIDIA A100. SGD is employed as the network optimizer with a momentum of 0.9 and an initial learning rate of 5e-5. The input resolution is set to 224 × 224, and the batch size is 64. For a warm-up, we first optimize the masked modeling objectives for 15,000 iterations and then jointly train all objectives end-to-end. After pre-training, we keep only the image and text encoders for downstream tasks. Using the CLIP (Radford et al., 2021) approach, zero-shot classification is done by matching the image and category name features from both encoders. Meanwhile, the image encoder can be used independently for fine-tuning classification tasks. About evaluation metrics, following previous works (Tiu et al., 2022; Xie et al., 2023), we adopt the area under the ROC curve (AUC), ACC, and F1 scores, which are standard metrics to evaluate classification tasks.
Table 1. Zero-shot performance of various pre-training models on three datasets.
Models CheXpert PadChest-Seen 8 Pneumonia
AUC F1 ACC AUC F1 ACC AUC F1 ACC
ConVIRT (Zhang et al., 2022) 52.10 35.61 57.43 74.31 23.58 80.12 79.21 55.67 75.08
GLORIA (Huang et al., 2021) 54.84 37.86 60.70 74.56 24.02 80.75 70.37 48.19 70.54
BioViL (Bannur et al., 2023) 60.01 42.10 66.13 71.17 20.75 79.58 84.12 54.49 74.43
BioViL-T (Bannur et al., 2023) 70.93 47.21 69.96 76.17 27.41 85.32 86.03 62.56 80.04
CheXzero (Tiu et al., 2022) 87.90 61.90 81.17 76.75 28.68 87.37 85.13 61.49 78.34
MedKlip (Wu et al., 2023b) 89.97 63.67 84.32 84.97 35.17 90.68 86.94 63.42 80.02
MMCLIP 90.53 67.05 86.04 85.79 33.50 91.19 86.16 62.57 80.46
+ Ensemble 91.45 69.18 86.28 86.31 35.61 92.16 86.95 63.30 80.51
5. Results
5.1. Zero-shot Classification
@@
We compare our MMCLIP with different pre-training methods, including ConVIRT (Zhang et al., 2022), GLORIA (Huang et al., 2021), BioViL (Bannur et al., 2023), BioViL-T(Bannur et al., 2023), CheXzero (Tiu et al., 2022), and MedKlip (Wu et al., 2023b), for zero-shot classification. The results in Table 1 and Table 2 show that our MMCLIP exhibits superior zero-shot performance across all these scenarios on three datasets, even for the unseen class recognition on the PadChest dataset. This may be because it is difficult for existing models to understand the feature information contained in the whole image and report, while the proposed AttMIM and EntMLM can utilize the rich multimodal information of the medical data itself to help our MMCLIP to understand the more critical medical knowledge, and even understand the complex medical features that have not been mentioned in the report of the training set.
@@
Specifically, in Table 1, compared to the second-best MedKlip, the AUC scores of our MMCLIP have increased as follows: from 89.97 to 90.53 on the CheXpert dataset, from 84.97 to 85.79 on the PadChest-seen dataset, and from 85.94 to ours 86.16 on the RSNA Pneumonia dataset. Meanwhile, we used the ensemble learning method to combine the weights obtained from 6 different groups of parameters, which can improve the performance further on most datasets. In Table 2, we comparatively analyze the zero-shot performance of MMCLIP on the PadChest dataset for unseen classes. Only 8 categories in PadChest strictly match the visible categories in the MIMIC data, so we only report the test results of 8 visible categories in Table 1. There are 12 categories that either cannot be strictly matched, but cannot be counted as unseen classes, or do not appear in the test set, so we report at most 173 unseen classes. Our analysis of the 10 and 20 most numerous of these unseen classes reveals that the zero-shot performance of MMCLIP has a 4% to 5% lift.
Table 2. Comparison of zero-shot AUC scores on the PadChest dataset. Unseen 10 represents the 10 categories with the highest number of unseen categories.
@@
CheXzero (Tiu et al., 2022) 59.75 63.11 65.76
MedKlip (Wu et al., 2023b) 60.07 53.73 64.16
MMCLIP 65.09 67.14 70.00
Figure 6. Performance comparison under different labeled data for fine-tuning. Unpaired data like PadChest can work with pair data to improve image encoder performance on the fine-tuning task.
5.2. Fine Tuning
Existing multi-modal frameworks are constrained as they require data that include both text and images, which limits the potential for the image encoder to leverage image-only data to improve its representation capabilities. To address this, we introduce a training approach designed to work effectively with unpaired, single-modal data. This new method allows the image encoder to better use image-only data, enhancing its representation power and overall performance. Our experiments involved two datasets, where we used 1%, 10%, and 100% of the data for fine-tuning, aligning with the methods in previous works (Wu et al., 2023b). As shown in Figure 6, our model significantly outperforms existing models in terms of AUC scores across all datasets. Meanwhile, the performance of the model was further improved by incorporating image data from PadChest, demonstrating that compatibility with unpaired data is vital for representation learning.
5.3. Ablation Study
@@
5.3.1. Improvements from AttMIM
As shown in Table 3, when we randomly mask 75% of the features according to the setting of MIM in MAE (He et al., 2022) and directly splice this module with CLIP and initialize it with the weights of MAE decoder, we can see that the model performance improves by 0.42% of AUC score when we train it directly. However, CLIP and MIM rely on different training strategies, with the latter tending to require smaller learning rates and longer training epochs. To harmonize the two, we first train the image encoder via the MIM reconstruction task for 15,000 iterations, as warm Up process. With the Warm Up, the model further improves the AUC by 1.1%. Meanwhile, we comparatively validate varying single strategies with mask ratios of 0.75, and the AttMIM strategy with the fusion of the three strategies. Experimental results show that three attention-based mask strategies can help the model understand key features better than random. As a fusion of the three strategies, AttMIM can more comprehensively improve related capabilities.
@@ 5.3.1. Improvements from AttMIM
Table 3. Zero-shot AUC on PadChest. ’Unseen 10’ refers to the 10 most frequently unseen categories.
@@
Table 4. Performance of different text features masking strategies on the new CheXpert validation set.
@@
5.3.2. Improvements from EntMLM
@@
Figure 7. AUC and F1 under different mask ratios of AttMIM.
Figure 8. AUC and F1 under different mask ratios of EntMLM.
6. Conclusion
In this paper, we propose MMCLIP, an advanced medical VLP framework that aims to overcome major challenges in medical image and text analysis by introducing an innovative cross-modal attention masking modeling mechanism. By fusing AttMIM and EntMLM, MMCLIP can accurately capture key pathological features in medical data and improve the model’s ability to understand complex medical scenes. In addition, we also use disease prompts to make the model compatible with unpaired data. Through pre-training on two large-scale medical datasets and validation on multiple downstream tasks, MMCLIP demonstrates its superior performance, not only achieving state-of-the-art results in the zero-shot setting but also in the fine-tuning setting. This result proves the practicality and effectiveness of MMCLIP in deep learning applications in the medical field.
Appendix
Appendix A Pre-training Dataset
MIMIC-CXR (Johnson et al., 2019) encompasses more than 227k investigations with paired imagery and report data, originating from 65,379 individuals across varied scanning instances. Each investigation might contain one or two images (diverse scanning perspectives), summing up to a total of 377,110 images.
PadChest (Bustos et al., 2020) is a comprehensive, chest x-ray dataset with 160k+ images from 67k patients, acquired at San Juan Hospital, Spain (2009-2017), inclusive of six varied positional views and rich associated patient/report data. Reports, categorized into 174 radiographic findings, 19 differential diagnoses, and 104 anatomical locations, utilize UMLS terminology and incorporate a hierarchical taxonomy. Limited by the fact that PadChest pairs are labelled in Spanish, this experiment only uses the image of this dataset.
Appendix B Datasets for Downstream Tasks
CheXpert (Irvin et al., 2019), is a large-scale dataset, containing 224,316 labelled chest X-rays from 65,240 patients. It covers 14 conditions including Fracture, Edema, Consolidation, Enlarged Cardiom, Cardiomegaly, Lung Lesion, Lung Opacity, Pneumonia, Atelectasis, Pneumothorax, Pleural Effusion, Pleural Other, Support Devices, and No Finding. This dataset is involved in zero-shot and linear probing in our downstream tasks, both using the officially released test set with 500 images, which is consistent with the existing work (Wu et al., 2023b; Xie et al., 2023).
COVIDx (Pavlova et al., 2022) datasets is used for COVID-19 diagnosis. COVIDx has 29,986 images from 16,648 patients, used for the binary classification task of COVID symptoms, with a 70%/20%/10% split for training, validation, and testing. This dataset is utilized for fine-tuning purposes, following the existing work (Wu et al., 2023b).
Pneumonia(Anouk Stein, 2018) comprises over 260,000 frontal view chest X-ray images along with their corresponding pneumonia opacity masks, as well as support binary classification for pneumonia symptom, which were collected by the Radiological Society of North America. Following the settings of previous work(Wu et al., 2023b), this dataset is divided into three parts: 60% for training, 20% for validation, and 20% for testing.
Xray14 (Wang et al., 2017) includes 112,120 frontal view X-ray images, each annotated with 14 disease labels derived from corresponding radiology reports. Following the methodology outlined in previous work(Wu et al., 2023b), the dataset is divided into three subsets: 70% of the data is allocated for training, 10% for validation, and 20% for testing. This division is carried out at the patient level to ensure that images from the same patient are not present across different subsets, thereby avoiding any overlap between the training, validation, and testing groups.
PadChest testset encompasses 193 different categories, which can be used to evaluate zero-shot classification performance for unseen classes. We first identify 20 categories that were present during the pretraining phase. Then, we pinpoint categories not seen during pretraining within the test set. Next, we rank these unseen categories based on their frequency and select groups of the top 10 and top 20 most frequent unseen categories, as well as a group including all 173 unseen categories. Performance tests for zero-shot classification are conducted separately on each of these groups.